OpenStudy (daniellelovee):

Statistics honors project check @ganeshie8

OpenStudy (daniellelovee):

hold on let me upload the work

ganeshie8 (ganeshie8):

1. Use the sample data on Pages 1 and 2 in the Worksheet.doc file to determine the proportion of times a flight in each sample was delayed. Remember, delay times are positive. Record the sample proportions on Page 4.

OpenStudy (daniellelovee):

yes those pages are from the worddocument I uploaded

ganeshie8 (ganeshie8):

Total how many observations are there in January 2011 ?

OpenStudy (daniellelovee):

40

ganeshie8 (ganeshie8):

In how many of those observations, the flights are delayed ? (positive delay)

OpenStudy (daniellelovee):

9 excluding the 0's because those are neutral right?

ganeshie8 (ganeshie8):

right, but im getting 10 count again

OpenStudy (daniellelovee):

yes sorry you are correct they are 10

ganeshie8 (ganeshie8):

So whats the proportion(ratio) of "number of delayed flights" to "toal number of flights" ?

OpenStudy (daniellelovee):

10/40

OpenStudy (daniellelovee):

1/4

ganeshie8 (ganeshie8):

Yep, similarly find the proportion for June 2011 table too

OpenStudy (daniellelovee):

alright brb

OpenStudy (daniellelovee):

3/40

ganeshie8 (ganeshie8):

Part 3 [After Lesson 6.6: Proportion Differences] Jan : \( p_1=1/4\) June : \(p_2=3/40\) Margin of error = \(Z^*\times\sqrt{\dfrac{p_1(1-p_1)}{n_1}+\dfrac{p_2(1-p_2)}{n_2}}\\~\\ =1.96\times\sqrt{\dfrac{\frac{1}{4}(1-\frac{1}{4})}{40}+\dfrac{\frac{3}{40}(1-\frac{3}{40})}{40}}\\~\\ =0.157\) Confidence interval = \((p_1-p_2) \pm \text{margin of error} \\~\\ =(\frac{1}{4}-\frac{3}{40}) \pm 0.157\\~\\ =(0.018,~0.332) \) Interpretation : We say that we are \(95\%\) confident that the difference between the two population proportions, \(p_1-p_2\), lies betweenhe calculated limits since, in repeated sampling, about \(95\%\) of the intervals constructed this way would contain \(p_1-p_2\)